I need help with this homework problem:
A swimming pool can be filled in 17 hours if water enters through a pipe alone, or in 20 hours if water enters through a hose alone. If water is entering through both the pipes and the hose, how long will it take to fill the pool?

Answer by jorel1380(3719)

1/17+1/20=1/t
20/340+17/340=1/t
37/340=1/t
37t=340
t=340/37
It takes 9.19 hours using both the pipes and the hose to fill the pool..